OpenStudy (anonymous):

csc2(π2−x)/1+tan2(x)

OpenStudy (alekos):

is that cosec squared, pi squared and tan squared

OpenStudy (alekos):

cosec^2(pi^2-x)/1+tan^2x ???

OpenStudy (anonymous):

yes sir

OpenStudy (alekos):

if you're trying to solve for x then it has to be equal to something

OpenStudy (alekos):

or are we simplifying?

OpenStudy (anonymous):

i need to Simplify the expression.

OpenStudy (alekos):

1+tan^2x = sec^2x

OpenStudy (alekos):

we can start with that

OpenStudy (alekos):

cosec^2(pi^2-x) = 1/sin^2(pi^2-x)

OpenStudy (alekos):

sec^2x =1/cos^2x

OpenStudy (alekos):

so we're left with cos^2x/sin^2(pi^2-x)

OpenStudy (alekos):

now are you sure it's pi squared at the start?

OpenStudy (anonymous):

pie/2 is what the problem says

OpenStudy (alekos):

that's what I thought

OpenStudy (anonymous):

will that change the problem

OpenStudy (anonymous):

can you use numerals to solve the problem